# Survey of Video Diffusion Models: Foundations, Implementations, and Applications

### v1

@@ 4.1.1 Image condition
Concatenation can also be done along the temporal axis (Shi et al., 2024; Ma et al., 2024b; Zhang et al., 2024i; Gu et al., 2024a; Dai et al., 2023), where the conditioning image is explicitly set as the first frame of the predictions. This method avoids the need for modifications to the model architecture.
@@
Moreover, with the incorporation of optical flow, image conditions can be utilized more explicitly. By warping the image features using optical flow and decoding the warped image, the generated frames effectively “copy” information from specific locations in the conditioning image. This approach is employed in several flow-based video generation methods (Ni et al., 2023; Li et al., 2024i).
@@ 4.1.2 Spatial condition
Boximator, SG-I2V, MotionBooth] ] ]
@@
Early examples include Tora (Zhang et al., 2024g), which enables users to input detailed trajectories for multiple objects, resulting in high-fidelity, multi-object videos that precisely follow the prescribed paths. Similarly, Motion-I2V (Shi et al., 2024) accepts sparse trajectory annotations and motion brushes directly on a reference image, ensuring temporal consistency and accurate object-level motion control.
@@
Scene-level conditions. While object-level conditions dictate the movement of individual elements, scene-level constraints shape the overall layout, geometry, and camera perspective. Methods like Streetscapes (Deng et al., 2024) rely on street maps, height maps, and camera trajectories to ensure spatial coherence and temporal stability in city-scale videos. SparseCtrl (Guo et al., 2024b) uses sparse sketches or depth maps at keyframes to propagate structural guidance through the sequence, reducing the need for dense inputs while maintaining spatial fidelity. CineMaster (Wang et al., 2025) further emphasizes 3D-aware scene manipulation by integrating precise 3D bounding box placement, flexible object and camera motion, and interactive layout control into a two-stage workflow.
@@ 4.1.3 Camera parameter condition
Figure 9: An example of integrating camera poses and object motion control for video generation (Fu et al., 2024).
@@
On the object motion side, various approaches link bounding-box–level or trajectory-based editing to camera control for deeper storytelling. Direct-A-Video (Yang et al., 2024i) allows users to specify camera movements such as panning or zooming alongside simple bounding-box manipulation for objects, removing the need for explicit motion annotations. MotionCtrl (Wang et al., 2024j) goes further by letting users provide detailed trajectory paths for both objects and the camera, integrating these into a diffusion-based model to achieve precise movements and strong frame-to-frame consistency. Likewise, 3DTrajMaster (Fu et al., 2024) incorporates 6DoF pose sequences for multi-entity 3D motion control, using a 3D-motion–grounded injector architecture that fuses text and pose embeddings and a gated attention mechanism within a diffusion pipeline. CineMaster (Wang et al., 2025) further couples bounding-box–based object manipulation with camera parameter specification in a two-stage pipeline. A Blender-based interface collects 3D bounding boxes and camera trajectories, which are automatically interpolated before guiding a specialized diffusion model that disentangles camera from object motion for enhanced 3D alignment. By merging camera parameters with object trajectories, these methods deliver more comprehensive control over how scenes unfold, enabling cohesive, lifelike video generation.
@@ 4.1.5 High-level video condition
Video Condition [Efficiency-Oriented Methods [Efficiency & Adaptation [SimDA, VidToMe,
@@
Tune-A-Video] ] ] [Specialized Control [Instruction-based
Editing [InstructVid2Vid, InsV2V] ] [Human-centric
Editing [Follow-Your-Pose, MagicPose,
@@
AniTalker, DPE] ] [Interactive Editing [DragVideo, Drag-A-Video, ReVideo,
VideoSwap] ] ] [Consistent Editing [Canonical Representation [LNA, VidEdit, DiffusionAtlas,
CoDeF, StableVideo, Lee et al.] ] [Latent Space
Manipulation [Text2Video-Zero, Rerender A Video,
@@
Diffusion Video Autoencoder, ColorDiffusor] ] [Structural & Motion
Guidance [Gen-1, MagicEdit, VideoComposer,
@@
VideoControlNet(Hu et al.)] ] [Attention-based
Editing [Tune-A-Video, Vid2Vid-Zero,
@@
AnyV2V, VidToMe, TokenFlow ] ] ] ]
@@
Consistent editing. A large body of work addresses how to preserve uniformity across frames, focusing on refined attention mechanisms, structural and motion guidance, latent space manipulation, and canonical representations. In terms of attention-based approaches, some of the earliest breakthroughs emerged when text-to-image diffusion models were extended to video. Tune-A-Video(Wu et al., 2023b) introduced sparse-causal spatio-temporal attention to adapt Stable Diffusion for one-shot video editing, while Vid2Vid-Zero(Wang et al., 2023d) developed a dense spatial-temporal attention module and cross-frame attention to capture longer-range temporal dependencies. Edit-A-Video(Shin et al., 2024) likewise addressed temporal flicker by injecting attention maps from previous frames and applying a novel temporal-consistent blending method. Further enhancements came from methods like EI2(Zhang et al., 2024j) and SimDA(Xing et al., 2024b), which minimized feature shifts via Shift-restricted Temporal Attention Modules, instance centering, spectral normalization, and Latent-Shift Attention to achieve smoother updates with fewer parameters. Some techniques also explicitly combine spatial and temporal features: UniEdit(Bai et al., 2024) and AnyV2V(Ku et al., 2024) fuse self-attention features across frames or integrate DDIM inversions to preserve both content and motion. Token-level operations further enhance frame alignment, with VidToMe(Li et al., 2024g) merging redundant tokens locally and globally, and TokenFlow(Geyer et al., 2023) identifying nearest-neighbor tokens to propagate consistent structure.
@@
Efficiency-oriented methods. Since diffusion-based pipelines can be computationally intensive, numerous works focus on streamlining editing without degrading consistency. SimDA(Xing et al., 2024b) employs Latent-Shift Attention and lightweight adapters, while VidToMe(Li et al., 2024g) uses Token Merging (ToMe) to compress attention tokens both locally and globally, saving compute cycles across frames. STEM Inversion(Li et al., 2024f) applies an EM algorithm to find compact bases for each frame, reducing the dimensionality of video representations and speeding up inference. On a similar note, Tune-A-Video(Wu et al., 2023b) demonstrates that carefully designed sparse spatio-temporal attention modules can maintain temporal stability while significantly cutting resource usage.
@@ 4.2.2 Video inpainting
Figure 11: Example of video inpainting with text instructions for video generation. The contents in the masked regions are inpainted under text instrubtions. Original figure from AVID (Zhang et al., 2024h)
While traditional methods typically adhere to a predefined problem setting with a specified masked area, recent diffusion methods aim to be more creative by incorporating language instructions and addressing more challenging cases through complex generation. One such example is shown in  Figure 11, where masks and instructions are provided to enable video editing applications as a video inpainting task (Zhang et al., 2024h).
@@
To better utilize the multi-modal capability of diffusion models, some research work has focused on adding inputs of other modalities, such as text instructions, into video inpainting, paving the way for real applications like text-guided video editing. AVID (Zhang et al., 2024h) builds on existing text-guided image-inpainting models and improves temporal consistency by integrating motion modules. They also proposed a structure guidance module tailored for different types of masks, and a middle-frame attention guidance method for longer videos. Subsequently, CoCoCo (Zi et al., 2024) proposes a more efficient motion capture module and an instance-aware region selection method for better “consistency, controllability, and compatibility” (and hence the name "CoCoCo").
@@ 4.2.3 Video interpolation and extrapolation/prediction
However, from the perspective of diffusion models, video interpolation and extrapolation can be tackled in a more unified framework. One example is illustrated in Figure 12. Therefore, we discuss both tasks together in this chapter.
@@
  • Difficulty: Traditional methods are typically designed for simple settings, where only one intermediate or future frame is predicted from context frames, with short time intervals between them to maintain smooth motion assumptions. In contrast, generative methods can handle more complex tasks, such as interpolating or predicting 5-10 frames at a time, and the context frames can be far apart in time. An extreme example would be having only the first and last frames of a 5-second video. In this case, generative models are well-suited to generate the entire video since most of the information is completely made up. Traditional methods, however, would struggle to handle this scenario due to lack of context. We will discuss more about long video generation in Section 4.5.
@@ 4.2.4 Video super-resolution
Video super-resolution is another well-established task in video restoration, focusing on enhancing low-resolution videos by upsampling the video frames to produce clearer and sharper high-resolution outputs. Early research focused more on image super-resolution algorithms, followed by video super-resolution methods, which introduced temporal consistency as an additional optimization objective. Some visual examples are shown in Figure 13.
@@
Some latest methods have also investigated different ways to combine super-resolution networks (mostaly a VAE) and diffusion networks (U-Net). SATeCo (Chen et al., 2024e) combined the super-resolution (VAE) and diffusion model (U-Net) by inserting the U-Net structure in between the VAE encoder and decoder, together with some adapation or alignment modules. The latest VEnhancer (He et al., 2024b) also shares a similar structure with additional space-time data augmentation and video-aware conditioning, which are trained to enhance AI-generated video on both space and time resolutions. In comparison, Upscale-A-Video (Zhou et al., 2024b) is more designed for long videos. They first utilize the U-Net to generate each video segment given optical flow inputs and perform a global recurrent latent propagation to ensure consistency across a longer time range. The VAE decoder is then applied to decode high-resolution images. Both structures have shown pros and cons on flexibility, fidelity and quality.
In addition to using diffusion models to enhance video resolutions, some research work has also studies how to enhance/adapt low-resolution image/video generation models so that it can also work well on high-resolution tasks.  Guo et al. (2025) propose a tuning-free method through a pivot replacement strategy that leverages reliable semantic guidance from the low-resolution model, as well as a tuning method that integrate learnable multi-scale upsampler modules to learn structural details at the new higher resolution. To follow up, FreeScale (Qiu et al., 2024c) introduces a tuning-free inference paradigm that fuses features from different perceptive scales and extracts the desired frequency for each part of the frame.
@@ 4.3 Personalization
By extending these techniques to scenarios without specialized video data, Still-Moving (Chefer et al., 2024) leverages text-to-image (T2I) customization and introduces lightweight spatial and temporal adapters for text-to-video pipelines. By training on duplicated still images (or “frozen videos”), it retains motion priors while aligning T2I and T2V features, enabling diverse personalization with only a handful of static images. In parallel, Dynamic Concepts (Abdal et al., 2025) addresses more inherently dynamic concepts—subjects whose characteristic motion is as integral to their identity as appearance—by fine-tuning a Diffusion Transformer in two stages. First, it learns an appearance “basis” from an unordered set of frames, and then refines temporal dynamics by freezing the learned parameters and focusing on motion consistency. Finally, Video Alchemist (Chen et al., 2025) expands personalization to multi-subject, open-set scenarios by fine-tuning a Diffusion Transformer, and personalizes both foreground objects and backgrounds while mitigating overfitting through targeted data augmentation.
@@ 4.4 Consistency
Diffusion models. Semantic consistency (Ren et al., 2024; Zhang et al., 2024i; Jiang et al., 2024b; Li et al., 2024g; Zhou et al., 2024d; Sun et al., 2023b) ensures the semantic meaning including objects and relationships between objects are smooth between different frames. The biggest problem is semantic drifting, which refers to the gradual shift or change in the meaning or context of visual content as a video progresses. StableVideo (Chai et al., 2023), as a pioneer work, proposes to propagate the representation to the next frame to ensure consistency. Similarly, TokenFlow (Geyer et al., 2023) preserves consistency by explicitly propagating diffusion features based on inter-frame correspondences using sampled tokens for video editing. GLOBER (Sun et al., 2023a) employs a novel adversarial loss along with conditioning the generation on the global features extracted by a video encoder for global consistency. Specifically, for face video generation, EMO (Tian et al., 2024) preserves visual/identity consistency by encoding a reference image/video/audio and cross-attention for video generation.
@@
Apart from these methods, another popular paradigm for consistency is controlling the noise used in diffusion models. PYoCo (Ge et al., 2023) is first proposed to employ a novel correlated noise model, which consists of a noise shared among all frames and an individual noise conditioned on the previous frames. Following this, a $\int$-noise (Chang et al., 2024) is proposed to warp noise following the flow and preserve the spatial Gaussian property, which is able to preserve the motion and semantic consistency at the same time.
One application of consistency is long video generation, which poses high requirements for motion and semantic consistency. long video generation (Ouyang et al., 2024c) is always achieved by ensuring a smooth transition and semantic consistency among different frames. StoryDiffusion (Zhou et al., 2024d) proposes to use a novel consistent self-attention and a semantic space temporal prediction for object consistency. Similarly, StreamingT2V (Henschel et al., 2024) employs a long-term memory block, a short-term memory block, and also a new blending approach for infinitely long video generation. The short-term block focuses on consistency between adjacent frames, while the long-term block forces the model to depend on the first several frames to avoid forgetting. Later, SEINE (Chen et al., 2024d) proposes to employ a random-mask model for generating transitions based on textual descriptions. Multiple randomly masked images can be seen as images from different scenes and they can provide an overall comprehensive angle to consistency and long video generation. On the other side, 3D scene mesh (Fridman et al., 2023) is also used for generating videos, which can be seen as another application of consistency. Specifically, the video frames and meshes are updated through several pre-trained models and the consistency between different frames is preserved through the iterated refinement of the mesh.

### v2

@@ 4.1.1 Image condition
Concatenation can also be done along the temporal axis (Shi et al., 2024; Ma et al., 2024b; Zhang et al., 2024j; Gu et al., 2024a; Dai et al., 2023), where the conditioning image is explicitly set as the first frame of the predictions. This method avoids the need for modifications to the model architecture.
@@
Moreover, with the incorporation of optical flow, image conditions can be utilized more explicitly. By warping the image features using optical flow and decoding the warped image, the generated frames effectively “copy” information from specific locations in the conditioning image. This approach is employed in several flow-based video generation methods (Ni et al., 2023; Li et al., 2024k).
@@ 4.1.2 Spatial condition
Spatial Condition
User-Driven Spatial Interactions
Peekaboo, DragNUWA, DragAnything
Multi-Modal Controls
MVideo, DreamVideo-2, MotionPrompting
Scene-Level Conditions
Streetscapes, SparseCtrl, CeneMaster
Object-Level Conditions
Tora, Motion-I2V, FreeTraj,ObjCtrl-2.5D,
Boximator, SG-I2V, MotionBooth
@@
Early examples include Tora (Zhang et al., 2024h), which enables users to input detailed trajectories for multiple objects, resulting in high-fidelity, multi-object videos that precisely follow the prescribed paths. Similarly, Motion-I2V (Shi et al., 2024) accepts sparse trajectory annotations and motion brushes directly on a reference image, ensuring temporal consistency and accurate object-level motion control.
@@
Scene-level conditions. While object-level conditions dictate the movement of individual elements, scene-level constraints shape the overall layout, geometry, and camera perspective. Methods like Streetscapes (Deng et al., 2024) rely on street maps, height maps, and camera trajectories to ensure spatial coherence and temporal stability in city-scale videos. SparseCtrl (Guo et al., 2024b) uses sparse sketches or depth maps at keyframes to propagate structural guidance through the sequence, reducing the need for dense inputs while maintaining spatial fidelity. CineMaster (Wang et al., 2025a) further emphasizes 3D-aware scene manipulation by integrating precise 3D bounding box placement, flexible object and camera motion, and interactive layout control into a two-stage workflow.
@@ 4.1.3 Camera parameter condition
Figure 14: An example of integrating camera poses and object motion control for video generation (Fu et al., 2024a).
@@
On the object motion side, various approaches link bounding-box–level or trajectory-based editing to camera control for deeper storytelling. Direct-A-Video (Yang et al., 2024i) allows users to specify camera movements such as panning or zooming alongside simple bounding-box manipulation for objects, removing the need for explicit motion annotations. MotionCtrl (Wang et al., 2024j) goes further by letting users provide detailed trajectory paths for both objects and the camera, integrating these into a diffusion-based model to achieve precise movements and strong frame-to-frame consistency. Likewise, 3DTrajMaster (Fu et al., 2024a) incorporates 6DoF pose sequences for multi-entity 3D motion control, using a 3D-motion–grounded injector architecture that fuses text and pose embeddings and a gated attention mechanism within a diffusion pipeline. CineMaster (Wang et al., 2025a) further couples bounding-box–based object manipulation with camera parameter specification in a two-stage pipeline. A Blender-based interface collects 3D bounding boxes and camera trajectories, which are automatically interpolated before guiding a specialized diffusion model that disentangles camera from object motion for enhanced 3D alignment. By merging camera parameters with object trajectories, these methods deliver more comprehensive control over how scenes unfold, enabling cohesive, lifelike video generation.
@@ 4.1.5 High-level video condition
High-level
Video Condition
Efficiency-Oriented Methods
Efficiency & Adaptation
SimDA, VidToMe,
@@
Tune-A-Video
Specialized Control
Instruction-based
Editing
InstructVid2Vid, InsV2V
Human-centric
Editing
Follow-Your-Pose, MagicPose,
@@
AniTalker, DPE
Interactive Editing
DragVideo, Drag-A-Video, ReVideo,
VideoSwap
Consistent Editing
Canonical Representation
LNA, VidEdit, DiffusionAtlas,
CoDeF, StableVideo, Lee et al.
Latent Space
Manipulation
Text2Video-Zero, Rerender A Video,
@@
Diffusion Video Autoencoder, ColorDiffusor
Structural & Motion
Guidance
Gen-1, MagicEdit, VideoComposer,
@@
VideoControlNet(Hu et al.)
Attention-based
Editing
Tune-A-Video, Vid2Vid-Zero,
@@
AnyV2V, VidToMe, TokenFlow
@@
Consistent editing. A large body of work addresses how to preserve uniformity across frames, focusing on refined attention mechanisms, structural and motion guidance, latent space manipulation, and canonical representations. In terms of attention-based approaches, some of the earliest breakthroughs emerged when text-to-image diffusion models were extended to video. Tune-A-Video(Wu et al., 2023b) introduced sparse-causal spatio-temporal attention to adapt Stable Diffusion for one-shot video editing, while Vid2Vid-Zero(Wang et al., 2023d) developed a dense spatial-temporal attention module and cross-frame attention to capture longer-range temporal dependencies. Edit-A-Video(Shin et al., 2024) likewise addressed temporal flicker by injecting attention maps from previous frames and applying a novel temporal-consistent blending method. Further enhancements came from methods like EI2(Zhang et al., 2024k) and SimDA(Xing et al., 2024b), which minimized feature shifts via Shift-restricted Temporal Attention Modules, instance centering, spectral normalization, and Latent-Shift Attention to achieve smoother updates with fewer parameters. Some techniques also explicitly combine spatial and temporal features: UniEdit(Bai et al., 2024) and AnyV2V(Ku et al., 2024) fuse self-attention features across frames or integrate DDIM inversions to preserve both content and motion. Token-level operations further enhance frame alignment, with VidToMe(Li et al., 2024h) merging redundant tokens locally and globally, and TokenFlow(Geyer et al., 2023) identifying nearest-neighbor tokens to propagate consistent structure.
@@
Efficiency-oriented methods. Since diffusion-based pipelines can be computationally intensive, numerous works focus on streamlining editing without degrading consistency. SimDA(Xing et al., 2024b) employs Latent-Shift Attention and lightweight adapters, while VidToMe(Li et al., 2024h) uses Token Merging (ToMe) to compress attention tokens both locally and globally, saving compute cycles across frames. STEM Inversion(Li et al., 2024g) applies an EM algorithm to find compact bases for each frame, reducing the dimensionality of video representations and speeding up inference. On a similar note, Tune-A-Video(Wu et al., 2023b) demonstrates that carefully designed sparse spatio-temporal attention modules can maintain temporal stability while significantly cutting resource usage.
@@ 4.2.2 Video inpainting
Figure 16: Example of video inpainting with text instructions for video generation. The contents in the masked regions are inpainted under text instrubtions. Original figure from AVID (Zhang et al., 2024i)
While traditional methods typically adhere to a predefined problem setting with a specified masked area, recent diffusion methods aim to be more creative by incorporating language instructions and addressing more challenging cases through complex generation. One such example is shown in  Figure˜16, where masks and instructions are provided to enable video editing applications as a video inpainting task (Zhang et al., 2024i).
@@
To better utilize the multi-modal capability of diffusion models, some research work has focused on adding inputs of other modalities, such as text instructions, into video inpainting, paving the way for real applications like text-guided video editing. AVID (Zhang et al., 2024i) builds on existing text-guided image-inpainting models and improves temporal consistency by integrating motion modules. They also proposed a structure guidance module tailored for different types of masks, and a middle-frame attention guidance method for longer videos. Subsequently, CoCoCo (Zi et al., 2024) proposes a more efficient motion capture module and an instance-aware region selection method for better “consistency, controllability, and compatibility” (and hence the name "CoCoCo").
@@ 4.2.3 Video interpolation and extrapolation/prediction
However, from the perspective of diffusion models, video interpolation and extrapolation can be tackled in a more unified framework. One example is illustrated in Figure˜17. Therefore, we discuss both tasks together in this chapter.
@@
  • Difficulty: Traditional methods are typically designed for simple settings, where only one intermediate or future frame is predicted from context frames, with short time intervals between them to maintain smooth motion assumptions. In contrast, generative methods can handle more complex tasks, such as interpolating or predicting 5-10 frames at a time, and the context frames can be far apart in time. An extreme example would be having only the first and last frames of a 5-second video. In this case, generative models are well-suited to generate the entire video since most of the information is completely made up. Traditional methods, however, would struggle to handle this scenario due to lack of context. We will discuss more about long video generation in Section˜4.5.
@@ 4.2.4 Video super-resolution
Video super-resolution is another well-established task in video restoration, focusing on enhancing low-resolution videos by upsampling the video frames to produce clearer and sharper high-resolution outputs. Early research focused more on image super-resolution algorithms, followed by video super-resolution methods, which introduced temporal consistency as an additional optimization objective. Some visual examples are shown in Figure˜18.
@@
Some latest methods have also investigated different ways to combine super-resolution networks (mostaly a VAE) and diffusion networks (U-Net). SATeCo (Chen et al., 2024f) combined the super-resolution (VAE) and diffusion model (U-Net) by inserting the U-Net structure in between the VAE encoder and decoder, together with some adapation or alignment modules. The latest VEnhancer (He et al., 2024b) also shares a similar structure with additional space-time data augmentation and video-aware conditioning, which are trained to enhance AI-generated video on both space and time resolutions. In comparison, Upscale-A-Video (Zhou et al., 2024b) is more designed for long videos. They first utilize the U-Net to generate each video segment given optical flow inputs and perform a global recurrent latent propagation to ensure consistency across a longer time range. The VAE decoder is then applied to decode high-resolution images. Both structures have shown pros and cons on flexibility, fidelity and quality.
In addition to using diffusion models to enhance video resolutions, some research work has also studies how to enhance/adapt low-resolution image/video generation models so that it can also work well on high-resolution tasks.  Guo et al. (2025a) propose a tuning-free method through a pivot replacement strategy that leverages reliable semantic guidance from the low-resolution model, as well as a tuning method that integrate learnable multi-scale upsampler modules to learn structural details at the new higher resolution. To follow up, FreeScale (Qiu et al., 2024c) introduces a tuning-free inference paradigm that fuses features from different perceptive scales and extracts the desired frequency for each part of the frame.
@@ 4.3 Personalization
By extending these techniques to scenarios without specialized video data, Still-Moving (Chefer et al., 2024) leverages text-to-image (T2I) customization and introduces lightweight spatial and temporal adapters for text-to-video pipelines. By training on duplicated still images (or “frozen videos”), it retains motion priors while aligning T2I and T2V features, enabling diverse personalization with only a handful of static images. In parallel, Dynamic Concepts (Abdal et al., 2025) addresses more inherently dynamic concepts—subjects whose characteristic motion is as integral to their identity as appearance—by fine-tuning a Diffusion Transformer in two stages. First, it learns an appearance “basis” from an unordered set of frames, and then refines temporal dynamics by freezing the learned parameters and focusing on motion consistency. Finally, Video Alchemist (Chen et al., 2025a) expands personalization to multi-subject, open-set scenarios by fine-tuning a Diffusion Transformer, and personalizes both foreground objects and backgrounds while mitigating overfitting through targeted data augmentation.
@@ 4.4 Consistency
Diffusion models. Semantic consistency (Ren et al., 2024; Zhang et al., 2024j; Jiang et al., 2024b; Li et al., 2024h; Zhou et al., 2024d; Sun et al., 2023b) ensures the semantic meaning including objects and relationships between objects are smooth between different frames. The biggest problem is semantic drifting, which refers to the gradual shift or change in the meaning or context of visual content as a video progresses. StableVideo (Chai et al., 2023), as a pioneer work, proposes to propagate the representation to the next frame to ensure consistency. Similarly, TokenFlow (Geyer et al., 2023) preserves consistency by explicitly propagating diffusion features based on inter-frame correspondences using sampled tokens for video editing. GLOBER (Sun et al., 2023a) employs a novel adversarial loss along with conditioning the generation on the global features extracted by a video encoder for global consistency. Specifically, for face video generation, EMO (Tian et al., 2024) preserves visual/identity consistency by encoding a reference image/video/audio and cross-attention for video generation.
@@
Apart from these methods, another popular paradigm for consistency is controlling the noise used in diffusion models. PYoCo (Ge et al., 2023) is first proposed to employ a novel correlated noise model, which consists of a noise shared among all frames and an individual noise conditioned on the previous frames. Following this, a $\int$-noise (Chang et al., 2024a) is proposed to warp noise following the flow and preserve the spatial Gaussian property, which is able to preserve the motion and semantic consistency at the same time.
One application of consistency is long video generation, which poses high requirements for motion and semantic consistency. long video generation (Ouyang et al., 2024c) is always achieved by ensuring a smooth transition and semantic consistency among different frames. StoryDiffusion (Zhou et al., 2024d) proposes to use a novel consistent self-attention and a semantic space temporal prediction for object consistency. Similarly, StreamingT2V (Henschel et al., 2024) employs a long-term memory block, a short-term memory block, and also a new blending approach for infinitely long video generation. The short-term block focuses on consistency between adjacent frames, while the long-term block forces the model to depend on the first several frames to avoid forgetting. Later, SEINE (Chen et al., 2024e) proposes to employ a random-mask model for generating transitions based on textual descriptions. Multiple randomly masked images can be seen as images from different scenes and they can provide an overall comprehensive angle to consistency and long video generation. On the other side, 3D scene mesh (Fridman et al., 2023) is also used for generating videos, which can be seen as another application of consistency. Specifically, the video frames and meshes are updated through several pre-trained models and the consistency between different frames is preserved through the iterated refinement of the mesh.